the length of a rectangle is 1 cm more than 4 times the width. the perimeter is 42 cm. find the length and width
let______=width
______=length
equation:
solve:

Answer by ewatrrr(24785)

Hi.
Let x be the width. then (4x +1) is the length.
.
Perimeter = 2*l + 2/w for a rectangle
.
question states***
2x * 2(4x+1) = 42cm
.
simplify and solve
2x + 8x + 2 = 42cm
.
10x = 40cm
.
x = 4cm (this is the width)
.
length = 17cm (4*4cm + 1)
.
check your answer
8cm + 34cm = 42cm